OpenStudy (anonymous):

Find the limit of the function by using direct substitution. lim/ x->0 (x^2-8) choices: 0 Does not exist -8 8

zepdrix (zepdrix):

Using direct substitution gives us: \[\Large\rm \lim_{x\to0}x^2-8=0^2-8\]Understand how that works?

OpenStudy (anonymous):

kind of.

OpenStudy (anonymous):

what happens to the 0^2

zepdrix (zepdrix):

0^2=0, yes? So 0 - 8 = ?

OpenStudy (anonymous):

-8

zepdrix (zepdrix):

yes, good job

OpenStudy (anonymous):

thxx